OpenStudy (anonymous):

Derek kicks a soccer ball off the ground and in the air, with an initial velocity of 31 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches? 14.2 feet 14.6 feet 15.0 feet 15.3 feet

OpenStudy (anonymous):

@iGreen , i got 15.0 but im not sure. can you explain this all out?

OpenStudy (igreen):

You have to find the vertex again.

OpenStudy (igreen):

Well to be more specific, you have to find the x-value of the vertex. You can use this formula: \(\dfrac{-b}{2a}\) Here, our function is in the form of \(ax^2 + bx + c\)

OpenStudy (igreen):

Therefore our points are: a = -16 b = 31 31 is from the velocity. \(\dfrac{-b}{2a}\) \(\dfrac{-31}{2(-16)}\) \(\dfrac{-31}{-32}\) Can you divide that?

OpenStudy (igreen):

Actually..we have to find the y-value of the vertex..lol.

OpenStudy (igreen):

Divide -31 / -32. @SouthernRebel101

OpenStudy (igreen):

@SouthernRebel101 Are you there?

OpenStudy (anonymous):

sorry @iGreen was away from computer people showed up at my door

OpenStudy (igreen):

Okay, can you divide -31 / -32?

OpenStudy (anonymous):

i got 0.96875

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yep, now we can plug this back in to find the y-value: H(t) = -16t^2 + 31t H(t) = -16(0.97)^2 + 31(0.97) H(t) = -16(0.9409) + 30.07 H(t) = -15.0544 + 30.07 Can you add that? @SouthernRebel101

OpenStudy (anonymous):

all of them added up?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Just add -15.0544 + 30.07

OpenStudy (anonymous):

i got 45.1244 @iGreen

OpenStudy (igreen):

That's -15.0544, not positive.

OpenStudy (anonymous):

oh, oops did not see that one seocnd

OpenStudy (anonymous):

15.0156, so was i right? C? @iGreen

OpenStudy (igreen):

Yep, C is right.

OpenStudy (anonymous):

:D YAY! thankyou so much

OpenStudy (igreen):

Np.